# On model emulation and closure tests for 3+1D relativistic heavy-ion collisions

### v2

@@ IV.4.3 Logarithmic transformation to training datasets
In this section, we discuss the impacts of training with the logarithm of the multiplicity observables on the accuracy of the GP emulations. Performing logarithmic transformation on particle multiplicities was found to be an effective trick for yielding more accurate GP emulators Nijs et al. (2021b); Everett et al. (2021a). It is suspected that $\ln(\rm{d}N/\rm{d}y)$ reduces the relative variations in the training dataset. However, the scaling transformation described in Sec. III should be a more systematic way to adjust the relative variations in the training data.
@@ IV.4.4 PCA Transformation to Functional Model Parameters
To explain 99% of the variance in the training design, we need three principal components for $\tilde{\eta}(\mu_{B})$ and $\langle y_{\rm loss}\rangle(y_{\mathrm{init}})$ and six principal components for $\tilde{\zeta}(T,\mu_{B})$. Since the functional parametrizations for $\tilde{\eta}(\mu_{B})$ and $\langle y_{\rm loss}\rangle(y_{\mathrm{init}})$ are simple piece-wise linear functions (see Eqs. (7) and (8)), the number of principal components needed is the same as the number of model parameters. For the QGP’s specific bulk viscosity, the parameterization in Eq. (9) is non-linear. The PCA transformation indicates that two more degrees of freedom are needed in addition to the four model parameters.
@@ IV.4.5 Active Learning Approach
We denote our approach here as “active learning” although we only have one iteration in our exercise. In principle, this technique can be applied in multiple iterations to increase the accuracy in the interesting regions of the high dimensional parameter space. This allows reducing the computational cost in a Bayesian parameter extraction using a computationally expensive physics model. A similar approach was employed in previous works Fan et al. (2023); Liyanage et al. (2023).
@@ V.1 MCMC Setups
Our analysis in this paper utilizes two different implementations of posterior sampling techniques. The first employs affine invariant MCMC, leveraging the emcee Python package Foreman-Mackey et al. (2013). Secondly, we employ Parallel Tempering Langevin Monte Carlo (PTLMC), implemented in the surmise package Plumlee et al. (2023). The parallel tempering part enhances the exploration of the target distribution by running multiple chains at different “temperatures” and exchanging states periodically, which scans the parameter space more accurately and reduces the probability of getting stuck in a local minimum. This technique is of advantage if the posterior distribution is a multi-modal distribution. Additionally, some Langevin dynamics are incorporated to explore complex posterior distributions more efficiently Roberts and Stramer (2002).
@@ V.2 Closure Test
To further quantify the closure test and compare results from different setups, introduced in Tab. 3 we will compare a metric $\Delta$ from the information-theoretic approach similar to Ref. Fan et al. (2023). The main contribution to the GP emulator information loss, which is mainly determined by the second moment of the posterior distribution, using
@@
where $\bm{\theta}_{\rm truth}$ is the closure test point, $|\bm{\theta}_{\rm max}-\bm{\theta}_{\rm min}|$ is the allowed prior range and $p(\bm{\theta})$ is the posterior distribution, where the closure test point is not included. Different from the definition in Ref. Fan et al. (2023), we are dividing by $N_{\rm param.}$, the number of model parameters, to normalize the values of $\Delta$ between zero and one. The smaller the value, the more peaked is the posterior distribution around the true value. The last column of Tab. 3 summarizes the values of $\Delta$ for the different MCMC runs.
@@
$\alpha_{\rm shadowing}$ $0.027_{-0.010}^{+0.011}$ $0.032_{-0.012}^{+0.012}$ $0.043_{-0.039}^{+0.067}$ $0.028_{-0.010}^{+0.011}$ $0.021$
$y_{{\rm loss},2}$ $1.514_{-0.019}^{+0.017}$ $1.511_{-0.022}^{+0.021}$ $1.464_{-0.114}^{+0.115}$ $1.514_{-0.019}^{+0.016}$ $1.511$
@@
$y_{{\rm loss},6}$ $2.366_{-0.037}^{+0.038}$ $2.361_{-0.045}^{+0.042}$ $2.216_{-0.234}^{+0.250}$ $2.368_{-0.035}^{+0.030}$ $2.338$
$\sigma_{y_{\rm loss}}$ $0.298_{-0.034}^{+0.021}$ $0.295_{-0.027}^{+0.024}$ $0.317_{-0.168}^{+0.182}$ $0.303_{-0.019}^{+0.018}$ $0.317$
@@
$\alpha_{\text{preFlow}}$ $0.081_{-0.011}^{+0.012}$ $0.082_{-0.013}^{+0.017}$ $0.155_{-0.100}^{+0.139}$ $0.081_{-0.010}^{+0.013}$ $0.086$
$\eta_{0}$ $0.101_{-0.007}^{+0.006}$ $0.100_{-0.010}^{+0.010}$ $0.086_{-0.035}^{+0.040}$ $0.101_{-0.006}^{+0.006}$ $0.098$
$\eta_{2}$ $0.073_{-0.008}^{+0.009}$ $0.077_{-0.012}^{+0.013}$ $0.130_{-0.099}^{+0.118}$ $0.073_{-0.007}^{+0.009}$ $0.073$
@@
$\zeta_{\rm max}$ $0.190_{-0.005}^{+0.004}$ $0.189_{-0.006}^{+0.006}$ $0.182_{-0.030}^{+0.016}$ $0.189_{-0.005}^{+0.005}$ $0.188$
$T_{\zeta,0}\;[{\rm GeV}]$ $0.228_{-0.003}^{+0.003}$ $0.228_{-0.003}^{+0.003}$ $0.226_{-0.029}^{+0.019}$ $0.228_{-0.003}^{+0.003}$ $0.228$
$\sigma_{\zeta,+}\;[{\rm GeV}]$ $0.077_{-0.005}^{+0.001}$ $0.075_{-0.003}^{+0.003}$ $0.075_{-0.024}^{+0.023}$ $0.077_{-0.005}^{+0.001}$ $0.082$
$\sigma_{\zeta,-}\;[{\rm GeV}]$ $0.041_{-0.002}^{+0.003}$ $0.041_{-0.002}^{+0.003}$ $0.042_{-0.021}^{+0.013}$ $0.041_{-0.002}^{+0.003}$ $0.041$
@@ VII Conclusion
The code developed for this work is open source and can be downloaded from Roch and Jahan (2024).

### v3

@@ IV.4.3 Logarithmic transformation to training datasets
In this section, we discuss the impacts of training with the logarithm of the multiplicity observables on the accuracy of the GP emulations. Performing logarithmic transformation on particle multiplicities was found to be an effective trick for yielding more accurate GP emulators Nijs et al. (2021b); Everett and others (2021a). It is suspected that $\ln(\rm{d}N/\rm{d}y)$ reduces the relative variations in the training dataset. However, the scaling transformation described in Sec. III should be a more systematic way to adjust the relative variations in the training data.
@@ IV.4.4 PCA Transformation to Functional Model Parameters
To explain 99% of the variance in the training design, we need three principal components for $\tilde{\eta}(\mu_{B})$ and $\langle y_{\rm loss}\rangle(y_{\mathrm{init}})$ and six principal components for $\tilde{\zeta}(T,\mu_{B})$. Since the functional parametrizations for $\tilde{\eta}(\mu_{B})$ and $\langle y_{\rm loss}\rangle(y_{\mathrm{init}})$ are simple piece-wise linear functions (see Eqs. (7) and (8)), the number of principal components needed is the same as the number of model parameters. For the QGP’s specific bulk viscosity, the parameterization in Eq. (9) is non-linear. The PCA transformation indicates that two more degrees of freedom are needed in addition to the four model parameters. We checked that the averaged relative RMS errors are less than 5% between the original parameterizations and the ones reconstructed using the truncated principal components.
@@ IV.4.5 Active Learning Approach
We denote our approach here as “active learning” although we only have one iteration in our exercise. In principle, this technique can be applied in multiple iterations to increase the accuracy in the interesting regions of the high dimensional parameter space. This allows reducing the computational cost in a Bayesian parameter extraction using a computationally expensive physics model. A similar approach was employed in previous works Fan and others (2023); Liyanage et al. (2023).
@@
Let us also mention here that including too many HPP points when training GP emulators can significantly reduce the emulator uncertainty in certain phase space regions. The non-uniform size of the emulator uncertainty over the entire parameter space could potentially introduce bias (small likelihood) to the parameter space near HPP when performing Bayesian inference with real experimental measurements Jahan et al. (2024). Based on the improvement in our closure test, we are confident that adding a small amount of HPP points ($\sim 10\%$ of the LHD points used) for the training does not introduce a noticeable bias in the Bayesian inference.
@@ V.1 MCMC Setups
Our analysis in this paper utilizes two different implementations of posterior sampling techniques. The first employs affine invariant MCMC, leveraging the emcee Python package Foreman_Mackey_2013. Secondly, we employ Parallel Tempering Langevin Monte Carlo (PTLMC), implemented in the surmise package surmise2023. The parallel tempering part enhances the exploration of the target distribution by running multiple chains at different “temperatures” and exchanging states periodically, which scans the parameter space more accurately and reduces the probability of getting stuck in a local minimum. This technique is of advantage if the posterior distribution is a multi-modal distribution and additionally converges faster than without parallel tempering. Additionally, some Langevin dynamics are incorporated to explore complex posterior distributions more efficiently Roberts2002-om.
@@ V.2 Closure Test
To further quantify the closure test and compare results from different setups, introduced in Tab. 3 we will compare a metric $\Delta$ from the information-theoretic approach similar to Ref. Fan and others (2023). The main contribution to the GP emulator information loss, which is mainly determined by the second moment of the posterior distribution, using
@@
where $\bm{\theta}_{\rm truth}$ is the closure test point, $|\bm{\theta}_{\rm max}-\bm{\theta}_{\rm min}|$ is the allowed prior range and $p(\bm{\theta})$ is the posterior distribution, where the closure test point is not included. Different from the definition in Ref. Fan and others (2023), we are dividing by $N_{\rm param.}$, the number of model parameters, to normalize the values of $\Delta$ between zero and one. The smaller the value, the more peaked is the posterior distribution around the true value. The last column of Tab. 3 summarizes the values of $\Delta$ for the different MCMC runs.
@@
$\alpha_{\rm shadowing}$ $027_{-0.010}^{+0.011} 0.0$$32_{-0.012}^{+0.012} 0.0$$43_{-0.039}^{+0.067} 0.0$$28_{-0.010}^{+0.011} 0.0$210.021
$y_{{\rm loss},2}$ $1.514_{-0.019}^{+0.017}$ $1.511_{-0.022}^{+0.021}$ $1.464_{-0.114}^{+0.115}$ $4_{-0.019}^{+0.016} 1.5111.5$11
@@
$y_{{\rm loss},6}$ $2.366_{-0.037}^{+0.038}$ $361_{-0.045}^{+0.042} 2.2$$216_{-0.234}^{+0.250}$ $2.368_{-0.035}^{+0.030}$ $2.338$
$\sigma_{y_{\rm loss}}$ $0.298_{-0.034}^{+0.021}$ $0.295_{-0.027}^{+0.024}$ $17_{-0.168}^{+0.182} 0.303$$03_{-0.019}^{+0.018}$ $0.317$
@@
$\alpha_{\text{preFlow}}$ $081_{-0.011}^{+0.012} 0.0$$082_{-0.013}^{+0.017}$ $155_{-0.100}^{+0.139} 0.0$$81_{-0.010}^{+0.013} 0.0$860.086
$\eta_{0}$ $0.101_{-0.007}^{+0.006}$ $100_{-0.010}^{+0.010} 0.0$$086_{-0.035}^{+0.040}$ $101_{-0.006}^{+0.006} 0.0$980.098
$\eta_{2}$ $3_{-0.008}^{+0.009} 0.0770.0$$7_{-0.012}^{+0.013}$ $130_{-0.099}^{+0.118} 0.0$$73_{-0.007}^{+0.009} 0.0$730.073
@@
$\zeta_{\rm max}$ $0.190_{-0.005}^{+0.004}$ $0.189_{-0.006}^{+0.006}$ $0.182_{-0.030}^{+0.016}$ $9_{-0.005}^{+0.005} 0.1880.1$88
$T_{\zeta,0}\;[{\rm GeV}]$ $0.228_{-0.003}^{+0.003}$ $0.228_{-0.003}^{+0.003}$ $0.226_{-0.029}^{+0.019}$ $_{-0.003}^{+0.003} 0.2280.228$
$\sigma_{\zeta,+}\;[{\rm GeV}]$ $077_{-0.005}^{+0.001} 0.0$$75_{-0.003}^{+0.003} 0.0$$5_{-0.024}^{+0.023} 0.077$$7_{-0.005}^{+0.001} 0.0$820.082
$\sigma_{\zeta,-}\;[{\rm GeV}]$ $041_{-0.002}^{+0.003} 0.0$$41_{-0.002}^{+0.003} 0.0$$42_{-0.021}^{+0.013} 0.0$$41_{-0.002}^{+0.003} 0.0$410.041
@@ VII Conclusion
It is worth mentioning, that in terms of the prediction reliability $\mathcal{H}$, the PCGP and PCSK emulators perform better than the Scikit GP, but for some observable points, there is a significant deviation from an honest prediction ($\mathcal{H}=0$) for all emulators currently used in the field. This is certainly something users of these methods should be aware of, and it leaves room for further studies on how this changes the values of extracted parameters.
@@
The code developed for this work is open source and can be downloaded from hendrik_roch_2024_11186662.